OpenStudy (anonymous):

What is the equation of the line which is parallel to the graph of y = −4x + 7 and which passes through the point (0, 3)? y = −4x + 3 y = 4x − 3 y = 5x − 11 y = 2x − 3 .i got the 1st one am i right?.. ;(

OpenStudy (anonymous):

im stuck between the 1st one and the 2cd onee...

OpenStudy (nikita2):

yes, you are right!

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

since your line is parellel to y=-4x +7 hence slope of your line will be -4; since your line is passing through (0,3) hence y-3 =-4(x-0) y=-4x +3

OpenStudy (anonymous):

thank you !